In each part below a function p and a function f are given. Treat p as a parent function and identify the horizontal graph trFor each part in the previous problem, sketch the graph of y = f(x) and y=p(x) on the same axes. Be sure to label your graphs (B)and(C)
Answer #1

bees (B.) peal = 5x solution. There compression and f(x) = Jan is a horizontal f(x) = 2x by ina factor of E ve - To plot y = 5x we put different values of x to get different values of y 3) put x = 1 we get y = 1) = put» put x=1 put X=2 put: x=3 put x=4 get y = 52 we get y = 2 we get y =56 we get y = 58. -Ques (0) Pra ya and f(x) = (x+3)2 Solection! There is a horizontal shift of 3 unit to the left in graph 2 f(x) = (x+3) * To p> put x = 1, y = (?=0 3 put x = 1, y = (2)2 = 9 5 put X=-1, y = 61*=D =) put x = -2, y = (-2)=19 To plot y = (x+3)? put diffe- = (x+3y- НА 1,16) че